OpenStudy (anonymous):

What is the sum? 5/3 + (-1/5)

OpenStudy (triciaal):

+a - becomes a - 5/3 - 1/5 lcd of 3 and 5 =15 (15)(5/3) - 15(1/3) with denominator 15 (25 -3)/15 =22/15